To find the equation of a regression line, y=ax+b, you need these formulas:

a=r sy/sx b=y-ax

A Data set has an r value of 0.553. If the standard deviation of the x-coordinates is 3.996, and the standard deviation of the y-coordinates is 6.203, what is the slope of the line to three decimal places?


1
Expert's answer
2020-10-19T17:52:00-0400

The slope (a) can be calculated as follows:


"a=r\dfrac{s_Y}{s_X}=0.553\cdot \dfrac{6.203}{3.996}\approx0.858"

The slope of the line is "0.858."